Question 5 1.8 s-ㅛ+음+ +고ew? Round your final answer to two decimal places. Question6 1.8 s-ur +1㎡,24+ ? Round your final answer to two decimal places. Question 7 1.81 O None of the answers are comect 1.81 Question 8 .. and kg
Answer #1

Question 5.

Solution: we know sum of n terms GP = a1( 1- rn) / (1-r)

The given series can be arranged as:

S = 30( 1/1.072 + 1/1.073 + ….1/1.0740)

= 30 (1/1.072) (1-1/1.0740) / (1-1/1.07) [ Here, a = r = 1/1.07 and n = 40]

= 30 * 0.0533

= 1.60

Question 6

Solution: We know sum of GP up to Infinity is = a1 / 1-r = 30[(1/1.072)/ (1-1/1.07) ]

= 400.53